OpenStudy (mayankdevnani):

Challenging Question :- ( NOT IN EXAMINATION POINT OF VIEW) You have to identify the mistake in this challenging proof. (Given Below) [ REMEMBER:- This proof is absolutely wrong. don't write this proof any more in examination.. It's just a challenge and use your brain to identify the mistake] Thank You. I hope all can answer this.

OpenStudy (mayankdevnani):

Proof of 4=5 \[\huge \bf 16-36=25-45\] \[\huge \bf 16-36+\frac{81}{4}=25-45+\frac{81}{4}\] \[\huge \bf (4-\frac{9}{2})^2=(5-\frac{9}{2})^2\] \[\huge \bf 4-\frac{9}{2}=5-\frac{9}{2}\] \[\huge \bf 4=5\]

OpenStudy (mayankdevnani):

Identify the mistake.

hartnn (hartnn):

mistake while taking square root on both sides

hartnn (hartnn):

the answer to \(\sqrt {x^2} = |x| \)

OpenStudy (anonymous):

Goku always saving the day...

hartnn (hartnn):

so infact , final answer will be \(|4-9/2|=|5-9/2| \\ |-1/2|=|-1/2| \\ 1/2 = 1/2\)

OpenStudy (anonymous):

This is true though:\[ \left| 4-\frac 92 \right| = \left|5-\frac 92\right| \]

OpenStudy (anonymous):

************************************************* This is irrelevant!!! ************************************************* \[ (4-\frac{9}{2})^2=(5-\frac{9}{2})^2\]\[ (4-\frac{9}{2})^2-(5-\frac{9}{2})^2=0\]\[ [(4-\frac{9}{2})+(5-\frac{9}{2})][(4-\frac{9}{2})+(5-\frac{9}{2})]=0\]\[ [9-2(\frac{9}{2})][(9-2(\frac{9}{2})]=0\]\[0\times0 = 0\]Lol! This is fun!

hartnn (hartnn):

lol, what a nice proof that 0 x 0 = 0

OpenStudy (anonymous):

Thanks LOL!

OpenStudy (mayankdevnani):

Goood.......going.........

OpenStudy (anonymous):

Is this a direct proof, or a "I'm going to assume it is true because I know it is true, oh look it's true!" proof?

OpenStudy (mayankdevnani):

thank you . this is true proof

OpenStudy (anonymous):

Ah-huh! I didn't know it turned out like that though :\